Train A and B are traveling in the same direction on parallel tracks. Train A is traveling at 80 miles per hour and train B is traveling at 90 miles per hour. Train A passes a station at 1:15 P.M. If train B passes the same station at 1:27 P.M. , at what time will train B catch up to train A?
Answer by mananth(16949)

Train A and B are traveling in the same direction on parallel tracks. Train A is traveling at 80 miles per hour and train B is traveling at 90 miles per hour. Train A passes a station at 1:15 P.M. If train B passes the same station at 1:27 P.M. , at what time will train B catch up to train A?
train A is at station at 1.15 pm
Train B is at 1.27 pm
in 12 minutes train A has traveled 16 miles
Let train B travel x miles from the station when it catches up with train A
time taken by A to travel the distance x-16 miles = x-16 / 80
time taken by B to travel x miles is x/90
Both these times are same x/90 = x-16 /80
80x= 90( x-16)
80x-90x=-1440
-10x= -1440
x= 144 miles.
144 / 90 is the time taken by B to travel from station when it catches up with A
= 1.6 hours = 1hour and 36 minutes
1.27pm + 1hour 36minutes = 3.03 pm
Train b catches up at 3.03 pm